# AffectGPT: A New Dataset, Model, and Benchmark for Emotion Understanding with Multimodal Large Language Models

### v1

@@ 5 Results and Discussion
Table 3 compares the performance of MER-Caption with existing datasets. For a fair comparison, we use the same model architecture and experimental settings and only change the training data. For general instruction datasets, we further conduct filtering experiments to remove samples without emotion-related content, emphasizing emotion-related subsets. The excellent performance of MER-Caption proves the limitations of current datasets in addressing MER. On the one hand, general instruction datasets pay insufficient attention to emotion-related tasks. On the other hand, emotional description datasets often suffer from inadequate dataset scales or insufficient annotation quality. Therefore, our dataset can serve as an important complement to existing datasets. In Table 4, compared to the results without filtering or with only low-level filtering, our two-level filtering leads to a performance improvement, further verifying the effectiveness of our filtering technique. See Appendix M for more details.
@@
Analysis of Input Impact.
User Study.
We conduct a user study to evaluate the quality of our proposed dataset. Since MERR-Fine and MERR-Coarse [15] share some samples with our dataset, we randomly select 20 overlapping samples. We then hire four expert annotators and present them with two descriptions for each sample: one from our dataset and one from the other datasets. The annotators are asked to watch the video and select the more accurate description. As shown in Table 7, our dataset provides more accurate descriptions than both the model-based MERR-Coarse and the human-filtered MERR-Fine, thereby validating the effectiveness of our proposed annotation strategy.
@@
This paper adopts Qwen2.5 as the default LLM. In Figure 4, we further explore the impact of different LLMs. Experimental results show that the performance difference brought by LLM is limited. These results verify that the superior performance of AffectGPT over the existing MLLMs does not come from LLM but from our proposed emotion description dataset and model.
Figure 4: LLM impact analysis.
@@ 6 Conclusion
Emotion plays an important role in human communication, conveying human intentions and deep thoughts. As Minsky [1] stated: The question is not whether intelligent machines can have any emotions, but whether machines can be intelligent without any emotions. The development of MER technology can enhance the human-computer interaction experience.
@@
This paper does not involve the collection of new data. The original data comes from the unlabeled part of MER2024 [33], with permission from the dataset owners. The annotation process does not involve hiring external annotators, and no ethical issues are associated with this process. Additionally, we restrict the use of this dataset under the license of CC BY-NC 4.0, requiring researchers to use our dataset responsibly. Therefore, no ethical concerns are raised in this paper.
@@ A.1 Emotion Dataset
Emotion datasets are the foundation for building MER systems [21, 39]. Most research has focused on building categorical datasets, where basic emotions are first defined, and annotators are asked to select the most likely one [40] or multiple [16] labels from basic emotions. However, emotions are often diverse [41] and can coexist [42], making it challenging for categorical datasets to fully capture these complex emotions.
To address this, recent studies have shifted from categorical datasets to descriptive datasets, as emotion descriptions provide greater flexibility and enable the description of complex emotions in natural language. To construct such datasets, [14] used a human-based annotation strategy to capture the environment, body movements, facial expressions, and other emotion-related cues. However, the high annotation cost limits the scalability of these datasets. With the development of MLLMs, [15] used a more cost-effective automatic annotation method, where MLLMs are used to extract emotion-related descriptions from audio, facial expressions, and visual objects. However, they lacked pre-experimentation on MLLM selection, relying on empirical model choices, leading to insufficient label quality. In this paper, we propose a solution to balance label quality and dataset size. By leveraging high-quality human-based datasets to guide description generation and sample filtering, we achieve a quality-assured automatic annotation process and ultimately construct MER-Caption.
@@ A.2 Emotion Models
Emotion models are closely related to the training corpus. For categorical datasets, researchers often build classifiers to map the human’s multimodal information to his emotion labels. Apart from choosing the architecture (such as CNN, RNN, or Transformers), most research focuses on how to align and fuse multimodal information. For example, [43] introduced a decomposition module to split features into modality-specific and modality-invariant representations. [44] aligned different modalities at the word level and then learned time-dependent cross-modal interactions. [45] proposed using cross-attention to align features in the latent space. More recently, [29] conducted a fair comparison of various fusion and alignment strategies, showing that temporal-preserving features do not always outperform time-compressed features, suggesting that MER may be more suitable to solve from a global perspective.
For descriptive datasets, due to their natural language style output, the framework needs to shift from traditional discriminative methods to generative methods. With the development of LLMs and MLLMs, researchers have started to build models based on them. For example, [46] used Vicuna as the language model, jointly training emotion labels and descriptions. [24] used the instruction-aware Q-Former module to learn the mapping between input images and emotional descriptions. [15] integrated different encoders to understand multimodal inputs and used LLaMA-2 as an LLM decoder. However, current models either only focus on unimodal information [46, 24] or leave all cross-modal interactions to the LLM [15], which is insufficient for solving MER tasks with multimodal characteristics. To this end, we introduce the AffectGPT model in this paper.
@@ Appendix B Implementation Details
Our choice of unimodal encoders is guided by previous research [29], using CLIP ViT-L [30] as the visual encoder and HUBERT-L [31] as the acoustic encoder. Given the remarkable performance of Qwen-2.5 [47], we choose it as the LLM. To ensure training efficiency, we only fine-turn an extra LoRA module (in the LLM), projector, and pre-fusion branch, while freezing the weights of the LLM and unimodal encoders (see Figure 3). This approach reduces GPU memory usage and speeds up training. Additionally, through preliminary experiments, we found that pre-training on other instruction datasets followed by a second-stage training on MER-Caption did not lead to performance improvements. The primary reason is the large scale of our dataset and the limited focus on MER in current instruction datasets. Therefore, we do not perform multi-stage training in our experiments. All models are implemented in PyTorch and conducted training and inference on 80GB NVIDIA Tesla A100 GPU. During training, we set the maximum number of epochs to 60, each epoch contains 5000 iterations, and the batch size of each iteration is 3. To optimize all trainable parameters, we use the AdamW optimizer [48] and set the learning rate to 1e-5. For more implementation details, please refer to the code provided in the supplementary materials.
@@ Appendix C Details about MLLMs
Otter [49] Video, Text https://github.com/Luodian/Otter
VideoChat [50] Video, Text https://github.com/OpenGVLab/Ask-Anything/tree/main/video_chat
VideoChat2 [51] Video, Text https://github.com/OpenGVLab/Ask-Anything/tree/main/video_chat2
Video-LLaVA [52] Video, Text https://github.com/PKU-YuanGroup/Video-LLaVA
Video-LLaMA [53] Video, Text https://github.com/DAMO-NLP-SG/Video-LLaMA
Video-ChatGPT [54] Video, Text https://github.com/mbzuai-oryx/Video-ChatGPT
LLaMA-VID [55] Video, Text https://github.com/dvlab-research/LLaMA-VID
mPLUG-Owl [56] Video, Text https://github.com/X-PLUG/mPLUG-Owl
Chat-UniVi [26] Video, Text https://github.com/PKU-YuanGroup/Chat-UniVi
SALMONN [25] Audio, Text https://github.com/bytedance/SALMONN
Qwen-Audio [57] Audio, Text https://github.com/QwenLM/Qwen-Audio
SECap [58] Audio, Text https://github.com/thuhcsi/SECap
OneLLM [59] Audio, Video, Text https://github.com/csuhan/OneLLM
PandaGPT [60] Audio, Video, Text https://github.com/yxuansu/PandaGPT
Emotion-LLaMA [15] Audio, Video, Text https://github.com/ZebangCheng/Emotion-LLaMA
@@ Appendix F Choice of Description Generation Strategy
Table 9: Preliminary experiments. We choose $\mbox{F}_{\mbox{s}}$ as the primary metric.
@@
SECap [58] 45.72±0.09 54.52±0.15 39.37±0.05
SALMONN [25] 47.96±0.04 50.20±0.04 45.92±0.04
Video-ChatGPT [61] 50.52±0.06 54.03±0.04 47.44±0.07
mPLUG-Owl [56] 52.73±0.13 54.54±0.13 51.04±0.13
Chat-UniVi [26] 53.08±0.01 53.68±0.00 52.50±0.02
@@ Appendix G Prompt for Clue Merge
Even when modality conflicts exist (i.e., the emotions conveyed by audio, video, and text are not the same. as shown in Figure 6), GPT-3.5 can provide reasonable responses, primarily due to its powerful reasoning ability.
@@ Appendix I Video Duration Distribution
MER-UniBench is a comprehensive evaluation benchmark covering three typical tasks in MER, including fine-grained emotion recognition, basic emotion recognition, and sentiment analysis. Different tasks involve different datasets, and we provide their statistical information in Table 10.
Table 10: Dataset statistics in MER-UniBench.
Dataset Chosen Set # Samples Label Description
@@
Basic Emotion MER2023 [22] MER-MULTI 411 most likely label among six candidates
@@
MER2024 [33] MER-SEMI 1,169 most likely label among six candidates
@@
IEMOCAP [34] Session5 1,241 most likely label among four candidates
@@
MELD [23] Test 2,610 most likely label among seven candidates
@@
Sentiment Analysis CMU-MOSI [35] Test 686 sentiment intensity, ranging from [-3, 3]
@@
CMU-MOSEI [36] Test 4,659 sentiment intensity, ranging from [-3, 3]
@@
CH-SIMS [37] Test 457 sentiment intensity, ranging from [-1, 1]
@@
CH-SIMS v2 [38] Test 1,034 sentiment intensity, ranging from [-1, 1]
@@
OV-MERD+ is our newly collected dataset, an extended version of the previous OV-MERD [12]. Unlike traditional datasets, which select a single label from basic emotions, OV-MERD is a fine-grained emotion dataset that allows each sample to have a variable number of emotions, using any emotion not restricted to predefined taxonomies. OV-MERD initially contains 332 samples, and we further expand its dataset size, obtaining OV-MERD+.
@@
MER2023 [22] and MER2024 [33] are widely used in Chinese MER research, with MER2024 being an extended version of MER2023. The original data in both datasets come from movies and TV shows. They use various techniques to segment video clips, ensuring that each clip has only one person, with their speech content being relatively complete. To ensure annotation quality, they hire multiple annotators, each selecting the most likely label from six candidate emotions: worry, happy, neutral, angry, surprised, and sad. The final label is determined through majority voting.
@@
IEMOCAP [34] is one of the most widely used emotion datasets. It contains five sessions, each with a male and a female actor in a laboratory environment. The dataset includes the following emotion labels: anger, happiness, sadness, neutral, excitement, frustration, fear, surprise, and others. Following previous research [62], we choose the last session for testing, and use the first four emotions, and merge surprise and happiness into happiness.
@@
MELD [23] is an extension of the text-centered EmotionLines dataset [63], adding audio and video content. The raw data is derived from the Friends TV series. The dataset has seven emotion labels, and each sample is assigned to one of the most likely labels: anger, joy, sadness, neutral, disgust, fear and surprise.
@@
CMU-MOSI [35] and CMU-MOSEI [36] consist of opinion videos collected from online platforms. CMU-MOSEI is an extended version of CMU-MOSI, with more samples and a wider range of topics. In these datasets, each sample is labeled with a sentiment intensity score ranging from -3 to +3, where -3 represents extremely negative emotion and +3 represents extremely positive emotion.
@@
CH-SIMS [37] and CH-SIMS v2 [38] differ from the English-centered CMU-MOSI and CMU-MOSEI by focusing on emotions within the Chinese culture. The original data comes from movies, TV series, and shows. Similar to CMU-MOSI and CMU-MOSEI, these datasets also annotate sentiment intensity, but with a different range $[-1,1]$, where -1 represents extremely negative emotion and +1 represents extremely positive emotion.
@@
Since there is no universal definition of the emotion wheel, we follow previous work [12] and use five emotion wheels in this paper.
@@
Figure 9: Emotion wheel. We use five emotion wheels, all of which are derived from previous research [12].

### v2

@@ 5 Results and Discussion
Effectiveness of MER-Caption.
Table 3 compares the performance of MER-Caption with existing datasets. For a fair comparison, we use the same model architecture and experimental settings and only change the training data. For general instruction datasets, we further conduct filtering experiments to remove samples without emotion-related content, emphasizing emotion-related subsets. Specifically, we use the prompt in Appendix E and extract emotion labels from each instruction-answer pair. Samples yielding empty emotion outputs are removed.
In Table 3, the excellent performance of MER-Caption proves the limitations of current datasets in addressing MER. On the one hand, general instruction datasets pay insufficient attention to emotion-related tasks. On the other hand, emotional description datasets often suffer from inadequate dataset scales or insufficient annotation quality. Therefore, our dataset can serve as an important complement to existing datasets. Meanwhile, for the general instruction datasets, the filtering approach is less effective on the LLaVA and VideoChat datasets. We hypothesize that the detailed descriptions in non-emotion subsets may also provide valuable cues for inferring emotional states in some scenarios.
Furthermore, we would like to acknowledge that MER-Caption+ may contain inaccurate descriptions due to the use of an automatic annotation strategy without manual checks. However, the experimental results in Table 3 show that MER-Caption+ achieves significantly better performance than the manually annotated MAFW dataset. The main reason is that humans tend to focus on major clues, which can easily lead to incomplete descriptions. These results confirm that, despite the lack of manual checks in MER-Caption+, we can still ensure the quality of the labels. In the future, we will investigate other post-filtering techniques to further improve MER-Caption+’s annotation quality.
@@
As shown in Table 4, compared to the results without filtering or with only low-level filtering, our two-level filtering leads to a performance improvement, further verifying the effectiveness of our filtering technique. These findings underscore that dataset quality is as critical as quantity, and fewer training samples do not necessarily lead to worse performance. Please see Appendix M for more details.
@@
Analysis of Input Impact.
User Study.
We conduct a user study to evaluate the quality of our proposed dataset. Since MERR-Fine and MERR-Coarse (Cheng et al., 2024) share some samples with our dataset, we randomly select 20 overlapping samples. We then hire four expert annotators and present them with two descriptions for each sample: one from our dataset and one from the other datasets. The annotators are asked to watch the video and select the more accurate description. As shown in Table 7, our dataset provides more accurate descriptions than both the model-based MERR-Coarse and the human-filtered MERR-Fine, thereby validating the effectiveness of our proposed annotation strategy.
@@
This paper adopts Qwen2.5 as the default LLM. In Figure 4(a), we further explore the impact of different LLMs. Experimental results show that the performance difference brought by LLM is limited. These results verify that the superior performance of AffectGPT over the existing MLLMs does not come from LLM but from our proposed emotion description dataset and model.
(a) LLM
(b) Audio Encoder
(c) Video Encoder
Figure 4: Ablation studies on LLMs, audio encoders, and video encoders.
Choice of Audio and Video Encoders.
In Figures 4(b) and 4(c), the choice of audio and video encoder has a minimal impact on performance. This underscores that AffectGPT’s exceptional performance is primarily driven by our proposed high-quality, large-scale dataset and effective framework, rather than the specific acoustic or visual encoders used. For audio encoders (Figure 4(b)), ImageBind exhibits slightly inferior performance compared to other audio encoders. This may be attributed to the fact that other audio encoders are predominantly utilized in audio content understanding tasks (e.g., ASR), where audio content plays a critical role in emotion recognition. Similarly, for video encoders (Figure 4(c)), CLIP_VIT marginally outperforms EVA_CLIP and DINOv2, aligning with findings from MERBench (Lian et al., 2024c), a unified benchmark for traditional categorical frameworks. These results suggest that insights derived from traditional categorical frameworks, such as encoder selection, may also be applicable to MLLM-based descriptive frameworks.
Role of LoRA in LLMs.
In Table 8, we count the increase in trainable parameters when using LoRA for the LLM branch. The first row represents the model without the LoRA module. Experimental results show that fine-tuning the LLM with LoRA improves performance compared to models without LoRA. However, increasing the rank for LoRA-based models does not yield significant performance gains and instead increases computational costs.
Table 8: Impact of rank value in LoRA. In this table, we count the increase in trainable parameters when using LoRA for the LLM branch. The first row represents the model without the LoRA module.
Rank # Increased Parameters MER-UniBench
0 73.30
8 20,185,088 74.65
16 40,370,176 74.77
32 80,740,352 74.92
@@ 6 Conclusion
Acknowledgments
This work is supported by the Excellent Youth Program of State Key Laboratory of Multimodal Artificial Intelligence Systems (MAIS2024311), the National Natural Science Foundation of China (62201572, 62322120, 61831022, 62276259, U21B2010, 62271083, 62306316, 62176165, 62206136, 62476146), the Stable Support Projects for Shenzhen Higher Education Institutions (20220718110918001), the Young Elite Scientists Sponsorship Program by CAST (2024QNRC001), and the University of Oulu& Research Council of Finland Profi 7 (352788).
@@ Acknowledgments
Emotion plays an important role in human communication, conveying human intentions and deep thoughts. As Minsky (Minsky, 1988) stated: The question is not whether intelligent machines can have any emotions, but whether machines can be intelligent without any emotions. The development of MER technology can enhance the human-computer interaction experience.
@@
This paper does not involve the collection of new data. The original data comes from the unlabeled part of MER2024 (Lian et al., 2024b), with permission from the dataset owners. The annotation process does not involve hiring external annotators, and no ethical issues are associated with this process. Additionally, we restrict the use of this dataset under the license of CC BY-NC 4.0, requiring researchers to use our dataset responsibly. Therefore, no ethical concerns are raised in this paper.
@@ A.1 Emotion Dataset
Emotion datasets are the foundation for building MER systems (Wang et al., 2022; Chen et al., 2023). Most research has focused on building categorical datasets, where basic emotions are first defined, and annotators are asked to select the most likely one (Goodfellow et al., 2013) or multiple (Li et al., 2017) labels from basic emotions. However, emotions are often diverse (Demszky et al., 2020) and can coexist (Du et al., 2014), making it challenging for categorical datasets to fully capture these complex emotions.
To address this, recent studies have shifted from categorical datasets to descriptive datasets, as emotion descriptions provide greater flexibility and enable the description of complex emotions in natural language. To construct such datasets, Liu et al. (2022a) used a human-based annotation strategy to capture the environment, body movements, facial expressions, and other emotion-related cues. However, the high annotation cost limits the scalability of these datasets. With the development of MLLMs, Cheng et al. (2024) used a more cost-effective automatic annotation method, where MLLMs are used to extract emotion-related descriptions from audio, facial expressions, and visual objects. However, they lacked pre-experimentation on MLLM selection, relying on empirical model choices, leading to insufficient label quality. In this paper, we propose a solution to balance label quality and dataset size. By leveraging high-quality human-based datasets to guide description generation and sample filtering, we achieve a quality-assured automatic annotation process and ultimately construct MER-Caption.
@@ A.2 Emotion Models
Emotion models are closely related to the training corpus. For categorical datasets, researchers often build classifiers to map multimodal human information to corresponding emotion labels. Apart from choosing the architecture (such as CNN, RNN, or Transformers), most research focuses on how to align and fuse multimodal information. For example, Hazarika et al. (2020) introduced a decomposition module to split features into modality-specific and modality-invariant representations. Gu et al. (2018) aligned different modalities at the word level and then learned time-dependent cross-modal interactions. Tsai et al. (2019) proposed using cross-attention to align features in the latent space. More recently, Lian et al. (2024c) conducted a fair comparison of various fusion and alignment strategies, showing that temporal-preserving features do not always outperform time-compressed features, suggesting that MER may be more suitable to solve from a global perspective.
For descriptive datasets, due to their natural language style output, the framework needs to shift from traditional discriminative methods to generative methods. With the development of LLMs and MLLMs, researchers have started to build models based on them. For example, Huang et al. (2024) used Vicuna as the language model, jointly training emotion labels and descriptions. Xie et al. (2024) used the instruction-aware Q-Former module to learn the mapping between input images and emotional descriptions. Cheng et al. (2024) integrated different encoders to understand multimodal inputs and used LLaMA-2 as an LLM decoder. However, current models either only focus on unimodal information (Huang et al., 2024; Xie et al., 2024) or leave all cross-modal interactions to the LLM (Cheng et al., 2024), which is insufficient for solving MER tasks with multimodal characteristics. To this end, we introduce the AffectGPT model in this paper.
@@ Appendix B Implementation Details
Our choice of unimodal encoders is guided by previous research (Lian et al., 2024c), using CLIP ViT-L (Radford et al., 2021) as the visual encoder and HUBERT-L (Hsu et al., 2021) as the acoustic encoder. Given the remarkable performance of Qwen-2.5 (Yang et al., 2024), we choose it as the LLM. To ensure training efficiency, we only fine-tune an extra LoRA module (in the LLM), projector, and pre-fusion branch, while freezing the weights of the LLM and unimodal encoders (see Figure 3). We default to setting the rank in the LoRA module to 16. This approach reduces GPU memory usage and speeds up training. Additionally, through preliminary experiments, we found that pre-training on other instruction datasets followed by a second-stage training on MER-Caption did not lead to performance improvements. The primary reason is the large scale of our dataset and the limited focus on MER in current instruction datasets. Therefore, we do not perform multi-stage training in our experiments. All models are implemented in PyTorch and conducted training and inference on 80GB NVIDIA Tesla A100 GPU. During training, we set the maximum number of epochs to 60, each epoch contains 5000 iterations, and the batch size of each iteration is 3. To optimize all trainable parameters, we use the AdamW optimizer and set the learning rate to 1e-5. For more implementation details, please refer to the code provided in https://github.com/zeroQiaoba/AffectGPT.
@@ Appendix C Details about MLLMs
Otter (Li et al., 2023a) Video, Text https://github.com/Luodian/Otter
VideoChat (Li et al., 2023c) Video, Text https://github.com/OpenGVLab/Ask-Anything/tree/main/video_chat
VideoChat2 (Li et al., 2024a) Video, Text https://github.com/OpenGVLab/Ask-Anything/tree/main/video_chat2
Video-LLaVA (Lin et al., 2024) Video, Text https://github.com/PKU-YuanGroup/Video-LLaVA
Video-LLaMA (Zhang et al., 2023) Video, Text https://github.com/DAMO-NLP-SG/Video-LLaMA
Video-ChatGPT (Maaz et al., 2024) Video, Text https://github.com/mbzuai-oryx/Video-ChatGPT
LLaMA-VID (Li et al., 2024b) Video, Text https://github.com/dvlab-research/LLaMA-VID
mPLUG-Owl (Ye et al., 2023) Video, Text https://github.com/X-PLUG/mPLUG-Owl
Chat-UniVi (Jin et al., 2024) Video, Text https://github.com/PKU-YuanGroup/Chat-UniVi
SALMONN (Tang et al., 2023) Audio, Text https://github.com/bytedance/SALMONN
Qwen-Audio (Chu et al., 2023) Audio, Text https://github.com/QwenLM/Qwen-Audio
SECap (Xu et al., 2024) Audio, Text https://github.com/thuhcsi/SECap
OneLLM (Han et al., 2024) Audio, Video, Text https://github.com/csuhan/OneLLM
PandaGPT (Su et al., 2023) Audio, Video, Text https://github.com/yxuansu/PandaGPT
Emotion-LLaMA (Cheng et al., 2024) Audio, Video, Text https://github.com/ZebangCheng/Emotion-LLaMA
@@ Appendix F Choice of Description Generation Strategy
We would like to clarify that, in this paper, we do not use combined results for model selection. Instead, we rely on the performance of individual models. Specifically, for VLLM, Chat-UniVi outperforms mPLUG-Owl and Video-ChatGPT; for ALLM, SALMONN outperforms SECap. As a result, we employ the combination of Chat-UniVi and SALMONN for description generation. The combination experiments are primarily designed to demonstrate that integrating multimodal cues can enhance performance. In future work, we will conduct additional experiments where combined results are used for model selection. For example, leveraging the combination of SALMONN and Chat-UniVi for description generation.
Table 10: Preliminary experiments. We choose $\mbox{F}_{\mbox{s}}$ as the primary metric, as this metric considers both accuracy and completeness.
@@
SECap (Xu et al., 2024) 45.72±0.09 54.52±0.15 39.37±0.05
SALMONN (Tang et al., 2023) 47.96±0.04 50.20±0.04 45.92±0.04
Video-ChatGPT (Maaz et al., 2024) 50.52±0.06 54.03±0.04 47.44±0.07
mPLUG-Owl (Ye et al., 2023) 52.73±0.13 54.54±0.13 51.04±0.13
Chat-UniVi (Jin et al., 2024) 53.08±0.01 53.68±0.00 52.50±0.02
@@ Appendix G Prompt for Clue Merge
Even when modality conflicts exist (i.e., the emotions conveyed by audio, video, and text are not the same, as shown in Figure 6), GPT-3.5 can provide reasonable responses, primarily due to its powerful reasoning ability.
@@ Appendix I Video Duration Distribution
MER-UniBench is a comprehensive evaluation benchmark covering three typical tasks in MER, including fine-grained emotion recognition, basic emotion recognition, and sentiment analysis. Different tasks involve different datasets, and we provide their statistical information in Table 11. In this paper, we intentionally focus on single-person videos, as this allows us to eliminate interference from other speakers and reduce task difficulty. Multi-person MER belongs to another research topic and will be addressed in our future work.
Table 11: Dataset statistics in MER-UniBench. All datasets in our study focus on single-person videos.
Dataset Chosen Set # Samples Label Description Data Source
@@
Basic Emotion MER2023 MER-MULTI 411 most likely label among six candidates movies, TV series
@@
MER2024 MER-SEMI 1,169 most likely label among six candidates movies, TV series
@@
IEMOCAP Session5 1,241 most likely label among four candidates actor’s performance
@@
MELD Test 2,610 most likely label among seven candidates ”Friends” TV series
@@
Sentiment Analysis CMU-MOSI Test 686 sentiment intensity, ranging from [-3, 3] opinion videos in YouTube
@@
CMU-MOSEI Test 4,659 sentiment intensity, ranging from [-3, 3] opinion videos in YouTube
@@
CH-SIMS Test 457 sentiment intensity, ranging from [-1, 1] movies, TV series, and shows
@@
CH-SIMS v2 Test 1,034 sentiment intensity, ranging from [-1, 1] movies, TV series, and shows
@@
OV-MERD+ is our newly collected dataset, an extended version of the previous OV-MERD (Lian et al., 2024a). Unlike traditional datasets, which select a single label from basic emotions, OV-MERD is a fine-grained emotion dataset that allows each sample to have a variable number of emotions, using any emotion not restricted to predefined taxonomies. OV-MERD initially contains 332 samples, and we further expand its dataset size, obtaining OV-MERD+.
@@
MER2023 (Lian et al., 2023) and MER2024 (Lian et al., 2024b) are widely used in Chinese MER research, with MER2024 being an extended version of MER2023. The original data in both datasets comes from movies and TV shows. They use various techniques to segment video clips, ensuring that each clip has only one person, with their speech content being relatively complete. To ensure annotation quality, they hire multiple annotators, each selecting the most likely label from six candidate emotions: worry, happy, neutral, angry, surprised, and sad. The final label is determined through majority voting.
@@
IEMOCAP (Busso et al., 2008) is one of the most widely used emotion datasets. It contains five sessions, each with a male and a female actor in a laboratory environment. The dataset includes the following emotion labels: anger, happiness, sadness, neutral, excitement, frustration, fear, surprise, and others. Following previous research (Poria et al., 2017), we choose the last session for testing, and use the first four emotions, and merge surprise and happiness into happiness.
@@
MELD (Poria et al., 2019) is an extension of the text-centered EmotionLines dataset (Hsu et al., 2018), adding audio and video content. The raw data is derived from the Friends TV series. The dataset has seven emotion labels, and each sample is assigned to one of the most likely labels: anger, joy, sadness, neutral, disgust, fear, and surprise.
@@
CMU-MOSI (Zadeh et al., 2017) and CMU-MOSEI (Zadeh et al., 2018) consist of opinion videos collected from online platforms. CMU-MOSEI is an extended version of CMU-MOSI, with more samples and a wider range of topics. In these datasets, each sample is labeled with a sentiment intensity score ranging from -3 to +3, where -3 represents extremely negative emotion and +3 represents extremely positive emotion.
@@
CH-SIMS (Yu et al., 2020) and CH-SIMS v2 (Liu et al., 2022b) differ from the English-centered CMU-MOSI and CMU-MOSEI by focusing on emotions within the Chinese culture. The original data comes from movies, TV series, and shows. Similar to CMU-MOSI and CMU-MOSEI, these datasets also annotate sentiment intensity, but with a different range $[-1,1]$, where -1 represents extremely negative emotion and +1 represents extremely positive emotion.
@@
Since there is no universal definition of the emotion wheel, we follow previous work (Lian et al., 2024a) and use five emotion wheels in this paper.
@@
Figure 9: Emotion wheel. We use five emotion wheels, all of which are derived from previous research (Lian et al., 2024a).